April 29, 2013  Engineering Optimization  gENOguide

To appear in Engineering Optimization Vol. 00, No. 00, Month 20XX, 1–17

GUIDE

Engineering Optimization L^{A}TEX style guide for authors (Style 2 + Chicago author-date reference style)

A.N. Author^{a}^{∗} and I.T. Consultant^{b}

aTaylor & Francis, 4 Park Square, Milton Park, Abingdon, UK; ^{b}Institut für Informatik, Albert-Ludwigs-Universität, Freiburg, Germany

(v4.1 released April 2013)

This guide is for authors who are preparing articles for the Taylor & Francis journal Engineering Optimization (gENO) using the L^{A}TEX document preparation system and the class file gENO2e.cls, which is available via the journal’s homepage on the Taylor & Francis website. Authors planning to submit articles in L^{A}TEX are advised to use gENO2e.cls as early as possible in the creation of their files.

Keywords: submission instructions; source file coding; environments; references citation; fonts; numbering (Please provide three to five keywords taken from terms used in your manuscript)

Index to information contained in this guide

1. Introduction 1.1. The gENO document class 1.2. Submission of L^{A}TEX articles

to the journal

2. Using the gENO class file

3. Additional features

3.1. Footnotes to article titles

and authors’ names 3.2. Abstracts 3.3. Lists 3.4. Landscape pages

4. Some guidelines for using standard features 4.1. Sections 4.2. Illustrations (figures) 4.3. Tables 4.4. Theorem-like environments 4.5. Typesetting mathematics

4.5.1. Displayed mathematics 4.5.2. Bold math italic symbols 4.5.3. Bold Greek 4.5.4. Upright Greek characters and the upright partial derivative sign 4.6. Acknowledgements 4.7. Notes 4.8. References

4.8.1. References cited in the

text 4.8.2. The list of references

4.9. Appendices 4.10. gENO macros

5. Example of a section heading including small caps, italic, and bold Greek such as κ

6. gENO journal style

6.1. Hyphens, n-rules, m-rules and

minus signs 6.2. References 6.3. Maths fonts

7. Troubleshooting

8. Fixes for coding problems

9. Obtaining the gENO2e class file 9.1 Via the Taylor & Francis

website 9.2 Via e-mail

∗Corresponding author. Email: latex.helpdesk@tandf.co.uk

1. Introduction

In order to assist authors in the process of preparing a manuscript for Engineering Optimization (gENO), the journal’s layout style has been implemented as a L^{A}TEX 2ε class file based on the article document class. A BibTEX style file is also provided to assist with the formatting of your references in a style appropriate to that of the journal.

Commands that differ from or are provided in addition to the standard L^{A}TEX 2ε interface are explained in this guide. The guide alone is not intended as a substitute for an appropriate L^{A}TEX 2ε manual.

The gENOguide.tex file can also be used as a template for composing an article for submission by cutting, pasting, inserting and deleting text as appropriate, using the

L^{A}TEX environments provided (e.g. \begin{equation}, \begin{enumerate}).

Please note that the index following the abstract in this guide is provided for information only. An index is not required in submitted articles.

1.1 The gENO document class

The gENO class file preserves the standard L^{A}TEX 2ε interface such that any document that can be produced using article.cls can also be produced using the gENO document class. However, the measure (the width of the text on a page) is slightly narrower than the default for article.cls, therefore line breaks will change and some long equations may need to be reformatted accordingly.

If your article is accepted for publication in the journal, it will be typeset in Monotype Times. As most authors do not own this font, it is likely that the page make-up will alter with the change of font. For this reason, please ignore details such as slightly long lines of text, page stretching, or figures falling out of synchronization with their citations in the text: these details will be dealt with by the typesetter. Similarly, it is unnecessary to spend time addressing warnings in the log file—if your .tex file compiles to produce a PDF file that correctly reflects how you wish your article to appear, such warnings will not prevent your source files being imported into the typesetter’s program.

1.2 Submission of L^{A}TEX articles to the journal

Manuscripts for possible publication in the journal should be submitted to the Editors for review as directed in the journal’s Instructions for Authors, which may be found at

http://www.tandf.co.uk/journals/authors/genoauth.asp.

Manuscripts created using L^{A}TEX should be converted to PDF format prior to submission. The L^{A}TEX source files and any graphics files will be required in addition to the final PDF version when final, revised versions of accepted manuscripts are submitted.

‘Open-source’ L^{A}TEX 2ε should be used in preference to proprietary systems such as TCILaTeX or Scientific WorkPlace; similarly, class files such as REVTeX4 that produce a document in the style of a different publisher and journal should not be used for preference.

Authors who wish to incorporate Encapsulated PostScript artwork directly in their articles can do so by using Tomas Rokicki’s EPSF macros (which are supplied with the DVIPS PostScript driver). See Section 3.4, which also demonstrates how to treat landscape pages. Please remember to supply any additional figure macros you use with your article in the preamble before \begin{document}. Authors should not attempt to use implementation-specific \specials directly.

Ensure that any author-defined macros are gathered together in the source file, just before the \begin{document} command.

Please note that if serious problems are encountered with the coding of an article (missing author-defined macros, for example), it may prove necessary to divert the article to conventional typesetting, i.e. it will be re-keyed.

2. Using the gENO class file

If the file gENO2e.cls is not already in the appropriate system directory for L^{A}TEX 2ε files, either arrange for it to be put there, or copy it to your working folder. In order to use the gENO document class, replace the command \documentclass{article} at the beginning of your document with the command \documentclass{gENO2e}.

The following document-class options should not be used with the gENO class file:

• 10pt, 11pt, 12pt—unavailable;

• oneside, twoside—not necessary, oneside is the default;

• leqno, titlepage—should not be used;

• onecolumn—not necessary as it is the default style.

The geometry package and commands associated with it should also not be used to adjust the page dimensions.

3. Additional features

3.1 Footnotes to article titles and authors’ names

The \thanks control sequence may be used to produce a footnote to either the title or authors’ names. Footnote symbols for this purpose should be used in the order:

† (coded as \dagger), ‡ (\ddagger), § (\S), ¶ (\P), ∥ (\|), †† (\dagger\dagger),
‡‡ (\ddagger\ddagger), §§ (\S\S), ¶¶ (\P\P), ∥∥ (\|\|).

Note that footnotes to the main text will automatically be assigned the superscript symbols 1, 2, 3,... by the class file, beginning afresh on each page.^{1}

The title, author(s) and affiliation(s) should be followed by the \maketitle command. If preparing an anonymized version for peer review, \maketitle may follow directly after the title in order to shield the authors’ identities from the reviewers.

3.2 Abstracts

At the beginning of your article, the title should be generated in the usual way using the \maketitle command. Immediately following the title you should include an abstract. The abstract should be enclosed within an abstract environment. For example, the titles for this guide were produced by the following source code:

\title{{\itshape Engineering Optimization} \LaTeX\ style guide
for authors (Style 2 + Chicago author-date reference style)}
\author{A.N. Author$^{a}$$^{\ast}$\thanks{$^\ast$Corresponding
author. Email: latex.helpdesk@tandf.co.uk \vspace{6pt}} and I.T.
Consultant$^{b}$\\\vspace{6pt}  $^{a}${\em{Taylor \& Francis,
4 Park Square, Milton Park, Abingdon, UK}}; $^{b}${\em{Institut

1These symbols will be changed to the style of the journal by the typesetter during preparation of your proofs. If preferred, the endnotes package may be used instead to set the notes in consecutive order at the end of your text, before the bibliography.

f\"{u}r Informatik, Albert-Ludwigs-Universit\"{a}t, Freiburg,
Germany}}\\\received{v4.1 released April 2013} }
\maketitle
\begin{abstract}
This guide is for authors who are preparing articles for the Taylor
\& Francis journal {\em Engineering Optimization} ({\it gENO}\,)
using the \LaTeX\ document preparation system and the class file
{\tt gENO2e.cls}, which is available via the journal's homepage
on the Taylor \& Francis website. Authors planning to submit
articles in \LaTeX\ are advised to use {\tt gENO2e.cls} as early
as possible in the creation of their files.
\end{abstract}

3.3 Lists

The gENO class file provides numbered and unnumbered lists using the enumerate environment and bulleted lists using the itemize environment.

The enumerated list numbers each list item with arabic numerals by default, e.g.

(1) first item

(2) second item

(3) third item

was produced by

\begin{enumerate}
  \item first item
  \item second item
  \item third item
\end{enumerate}

Alternative numbering styles can be achieved by inserting an optional argument in square brackets to each \item, e.g. \item[(i)] first item to create a list numbered with roman numerals.

Unnumbered lists are also provided using the enumerate environment, e.g.

First unnumbered indented item without label.

Second unnumbered item. Third unnumbered item.

was produced by:

\begin{enumerate}
  \item[] First unnumbered indented item without label.
  \item[] Second unnumbered item.
  \item[] Third unnumbered item.
\end{enumerate}

Bulleted lists are provided using the itemize environment. For example,

• First bulleted item

• Second bulleted item

• Third bulleted item

was produced by:

\begin{itemize}
  \item First bulleted item
  \item Second bulleted item
  \item Third bulleted item
\end{itemize}

3.4 Landscape pages

If a table or illustration is too wide to fit the standard measure, it must be turned, with its caption, through 90^{∘} anticlockwise. Landscape illustrations and/or tables can be produced directly using the gENO2e style file using \usepackage{rotating} after \documentclass{gENO2e}. The following commands can be used to produce such pages.

\setcounter{figure}{2}
\begin{sidewaysfigure}
\centerline{\epsfbox{fig1.eps}}
\caption{This is an example of a figure caption.}
\label{landfig}
\end{sidewaysfigure}
\setcounter{table}{0}
\begin{sidewaystable}
  \tbl{The Largest Optical Telescopes.}
    \begin{tabular}{@{}llllcll}
    .
    .
    .
  \end{tabular}\label{tab1}
\end{sidewaystable}

Before any float environment, use the \setcounter command as above to fix the numbering of the caption. Subsequent captions will then be automatically renumbered accordingly.

4. Some guidelines for using standard features

The following notes are intended to help you achieve the best effects with the gENO2e class file.

4.1 Sections

L^{A}TEX 2ε provides five levels of section headings and they are all defined in the gENO2e class file:

(A) \section
(B) \subsection
(C) \subsubsection
(D) \paragraph
(E) \subparagraph

Numbering is automatically generated for section, subsection, subsubsection and paragraph headings. If you need additional text styles in the headings, see the examples in Section 5.

(a) An example of an individual figure sub-caption.

(b) An example of an individual figure sub-caption.

Figure 1. Example of a two-part figure with individual sub-captions showing that captions are flush left and justified if greater than one line of text, otherwise centred under the figure.

4.2 Illustrations (figures)

The gENO class file will cope with most positioning of your illustrations and you should not normally use the optional positional qualifiers of the figure environment, which would override these decisions. See ‘Instructions for Authors’ in the journal’s homepage on the Taylor & Francis website for how to submit artwork (please note that requests to supply figures separately from text are for the benefit of authors using Microsoft Word; authors using L^{A}TEX may include illustrations at the appropriate locations in their PDF file). The original source files of any illustrations will be required when the final, revised version is submitted. Authors should ensure that their figures are suitable (in terms of lettering size, etc.) for the reductions they intend.

Figure captions should appear below the figure itself, therefore the \caption command should appear after the figure. For example, Figure 1 with caption and sub-captions is produced using the following commands:

\begin{figure}
\begin{center}
\begin{minipage}{120mm}
\subfigure[An example of an individual figure sub-caption.]{
\resizebox*{6cm}{!}{\includegraphics{senu_gr1.eps}}}
\subfigure[An example of an individual figure sub-caption.]{
\resizebox*{6cm}{!}{\includegraphics{senu_gr2.eps}}}
\caption{\label{fig2} Example of a two-part figure with individual
 sub-captions showing that captions are flush left and justified if
 greater than one line of text, otherwise centred under the figure.}
\label{sample-figure}
\end{minipage}
\end{center}
\end{figure}
  The control sequences \epsfig{}, \subfigure{} and \includegraphics{} require

epsfig.sty, subfigure.sty and graphicx.sty. These are called by the gENO2e class file and are included with the L^{A}TEX package for this journal for convenience.

To ensure that figures are correctly numbered automatically, the \label{} command should be inserted just after \caption{}.

Table 1. Radio-band beaming model parameters for FSRQs and BL Lacs.

Class^{a}  γ_{1} γ_{2}^{b} ⟨γ⟩  G  |f|  θ_{c}
BL Lacs 5  36  7 −4.0 1.0 × 10^{−2} 10^{∘}
FSRQs  5  40 11 −2.3 0.5 × 10^{−2} 14^{∘}

aThis footnote shows what footnote symbols to use. bThis footnote shows the text turning over when a long footnote is added.

4.3 Tables

The gENO class file will cope with most positioning of your tables and you should not normally use the optional positional qualifiers of the table environment, which would override these decisions. The table caption appears above the body of the table in gENO style, therefore the \tbl command should appear before the body of the table.

The tabular environment can be used as illustrated here to produce tables with single thick and thin horizontal rules, which are allowed, if desired. Thick rules should be used at the head and foot only and thin rules elsewhere.

Commands to redefine quantities such as \arraystretch should be omitted. For example, Table 1 is produced using the following commands. Note that \rm will produce a roman character in math mode. There are also \bf and \it, which produce bold face and text italic in math mode.

\begin{table}
\tbl{Radio-band beaming model parameters for {FSRQs and BL Lacs.}}
{\begin{tabular}{@{}lcccccc}
\toprule
  Class$^{\rm a}$ & $\gamma _1$ & $\gamma _2$$^{\rm b}$
         & $\langle \gamma \rangle$
         & $G$ & $|{\bm f}|$ & $\theta _{c}$ \\
\colrule
  BL Lacs &5 & 36 & 7 & $-4.0$ & $1.0\times 10^{-2}$ & 10$^\circ$ \\
  FSRQs & 5 & 40 & 11 & $-2.3$ & $0.5\times 10^{-2}$ & 14$^\circ$ \\
\botrule
\end{tabular}}
\tabnote{$^{\rm a}$This footnote shows what footnote symbols to use.}
\tabnote{$^{\rm b}$This footnote shows the text turning over when a
 long footnote is added.}
\label{symbols}
\end{table}

To ensure that tables are correctly numbered automatically, the \label{} command should be inserted just before \end{table}.

Tables produced using the booktabs package of macros for typesetting tables are also compatible with the gENO class file.

4.4 Theorem-like environments

The gENO style provides for the following theorem-like environments.

Theorem 4.1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Lemma 4.2 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Corollary 4.3 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Proposition 4.4 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Definition 4.5 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Proof More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT. ■

Remark 1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT problem.

Algorithm 1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT problem.

These were produced by:

\begin{theorem}
More recent algorithms for solving the semidefinite programming
relaxation are particularly efficient, because they explore the
structure of the MAX-CUT.
\end{theorem}
\begin{lemma}
...
...
\end{lemma}
\begin{corollary}
...
...
\end{corollary}
\begin{proposition}
...
...
\end{proposition}
\begin{definition}
...
...
\end{definition}
\begin{proof}
...
...
\end{proof}
\begin{remark}
...

...
\end{remark}
\begin{algorithm}
...
...
\end{algorithm}

4.5 Typesetting mathematics

4.5.1 Displayed mathematics

The gENO class file will set displayed mathematics centred on the measure without equation numbers, provided that you use the L^{A}TEX 2ε standard control sequences open (\[) and close (\]) square brackets as delimiters. The equation

∑_{i=1}^{p} λ_{i} = trace(S) i ∈ ℝ

was typeset using the commands

\[
  \sum_{i=1}^p \lambda_i = {\rm trace}({\textrm{\bf S}})\qquad
  i\in {\mathbb R}
\].

For those of your equations that you wish to be automatically numbered sequentially throughout the text, use the equation environment, e.g.

∑_{i=1}^{p} λ_{i} = trace(S)  i ∈ ℝ  (1)

was typeset using the commands

\begin{equation}
  \sum_{i=1}^p \lambda_i = {\rm trace}({\textrm{\bf S}})quad
  i\in {\mathbb R}
\end{equation}

Part numbers for sets of equations may be generated using the subequations environment, e.g.

ερw_{tt}(s, t) = N[w_{s}(s, t), w_{st}(s, t)]_{s},  (2a)

w_{tt}(1, t) + N[w_{s}(1, t), w_{st}(1, t)] = 0,  (2b)

which was generated using the control sequences

\begin{subequations} \label{subeqnexample}
\begin{equation}

        \varepsilon \rho w_{tt}(s,t)
        =
        N[w_{s}(s,t),w_{st}(s,t)]_{s},
        \label{subeqnpart}
\end{equation}
\begin{equation}
        w_{tt}(1,t)+N[w_{s}(1,t),w_{st}(1,t)] = 0,
\end{equation}
\end{subequations}

This is made possible by the package subeqn, which is called by the class file. If you put the \label{} just after the \begin{subequations} line, references will be to the collection of equations, ‘(2)’ in the example above. Or, like the example code above, you can reference each equation individually—e.g. ‘(2a)’.

4.5.2 Bold math italic symbols

To get bold math italic you can use \bm, which works for all sizes, e.g.

\sffamily
\begin{equation}
   {\rm d}({\bm s_{t_{\bm u}}) = \langle{\bm\alpha({\sf{\textbf L}})}
   [RM({\bm X}_y + {\bm s}_t) - RM({\bm x}_y)]^2 \rangle
\end{equation}
\normalfont

produces

d(s_{t_{u}}) = ⟨α(L)[RM(X_{y} + s_{t}) − RM(x_{y})]^{2}⟩  (3)

Note that subscript, superscript, subscript to subscript, etc. sizes will take care of themselves and are italic, not bold, unless coded individually. \bm produces the same effect as \boldmath. \sffamily...\normalfont allows upright sans serif fonts to be created in math mode by using the control sequence ‘\sf’.

4.5.3 Bold Greek

Bold lowercase as well as uppercase Greek characters can be obtained by {\bm \gamma}, which gives γ, and {\bm \Gamma}, which gives Γ.

4.5.4 Upright lowercase Greek characters and the upright partial derivative sign

Upright lowercase Greek characters can be obtained with the class file by inserting the letter ‘u’ in the control code for the character, e.g. \umu and \upi produce μ (used, for example, in the symbol for the unit microns—μm) and π (the ratio of the circumference to the diameter of a circle). Similarly, the control code for the upright partial derivative

∂ is \upartial.

4.6 Acknowledgements

This unnumbered section, e.g. \section*{Acknowledgement(s)}, should be used for thanks, grant details, etc. and placed before any Notes or References sections.

4.7 Notes

This unnumbered section, e.g. \section*{Note(s)}, may be placed before any References section.

4.8 References

4.8.1 References cited in the text

References should be cited in the text in ‘Chicago author-date’ style, separated by semi-colons, e.g. (Green 2002; Harlow 1983; Wegener and Petty 1994) or ‘... see Smith (1985, 75)’. If there are two authors with the same surname, use the first initials with the surnames, e.g. (E. Johnson 2001; L. Johnson 1998). If there are three authors, list them all in every citation, e.g. (Kernis, Cornell, and Harlow 1993). For more than three authors, cite the first author’s name followed by et al. For two or more sources by the same author(s) in the same year, use lower-case letters (a, b, c,...) with the year to order the entries in the Reference list and use these letters with the year in the in-text citations, e.g. (Green 1981a,b). Further details on this reference style can be found in the journal’s Instructions for Authors.

Each bibliographical entry has a key, which is assigned by the author and used to refer to that entry in the text. In this document, the key ev94 in the citation form \citep{ev94} produces ‘(Evans 1994)’, and the keys in the citation \citep{Eri1984,GloRib51,PeaEtAl76} produce ‘(Ericsson and Simon 1984; Glover, Smith, and Ribeiro 1951; Pierce et al. 1976)’. The appropriate citation style for different situations can be obtained, for example, by \citet{hk97} for ‘Kern (1997)’, \citet{Bar91,Mil93,Maz91} for ‘Barlow (1991); Miller (1993); Mazzeo (1991)’, or \citealt{Holl04} for ‘Holland 2004’. Optional notes may be included at the beginning and end of a citation by the use of square brackets, e.g. \citep[see][123]{Agu95} produces ‘(see Agutter 1995, 123)’. Citation of the year alone may be produced by

\citeyear{cwm73}, i.e. ‘1973’, or \citeyearpar{glov86}, i.e. ‘(1986)’.

4.8.2 The list of references

References should be listed at the end of the main text in alphabetical order, then chronologically (earliest first), with full page ranges (where appropriate) and issue numbers (for journals paginated by issue). If a reference has more than ten named authors, list only the first seven, followed by ‘et al.’.

The following list shows some references prepared in the style of the journal:

References

Agutter, A. J. 1995. “The Linguistic Significance of Current British Slang.” Unpublished doctoral

dissertation, Edinburgh University, UK. Barlow, D. H. 1991. “Diagnoses, Dimensions, and DSM-IV [Special issue].” Journal of Abnormal

Psychology 100 (3). Ericsson, K. Anders, and Herbert A. Simon. 1984. Protocol Analysis: Verbal Reports as Data.

Cambridge, MA: MIT Press. Evans, W. A. 1994. “Approaches to Intelligent Information Retrieval.” Information Processing

and Management 7 (2): 147–168. Glover, Fred. 1986. “Hilbert Modular Forms and the Galois Representations Associated to

Hilbert–Blumenthal Abelian Varieties.” Ph.D. thesis, Harvard University, Cambridge, MA. Glover, Fred, Raymond Smith, and C. C. Ribeiro, eds. 1951. Lessons of the British War Economy.

vol. 5. 2nd ed. Westport, CT: Greenwood Press.

Holland, M. 2004. “Guide to Citing Internet Sources.” Accessed November 4, 2012.

http://www.bournemouth.ac.uk/library/using/guide to citing internet sourc.html. Kern, Hudson. 1997. “The Resurgent Japanese Economy and a Japan–United States Free Trade

Agreement.” In 4th International Conference on the Restructuring of the Economic and Political System in Japan and Europe, edited by C. Lambert and G. Holst. Milan, Italy, 21–25 May 1996. 147–156. Singapore: World Scientific.

Mazzeo, J. 1991. Comparability of Computer and Paper-and-Pencil Scores (College Board Rep.

No. 91). Princeton, NJ: Educational Testing Service. Miller, M. E. 1993. The Interactive Tester (Version 4.0) [Computer software]. Westminster, CA:

Psytek Services. Misner, Charles W. 1973. “Efficient Algorithms for Layer Assignment Problems.” In Gravitation

in a Collapsing Universe, Vol. 5 of Einstein’s Legacy, edited by Ignatius Gottlob. 2nd ed. 63–83.

San Francisco, CA: Freeman. Pierce, Ian F., Fred Glover, A. T. Ogilvy, and A. R. S. Clements. 1976. “A Model of Output,

Employment, Wages and Prices in the UK.” In Computing Tools for Modeling, Optimization and Simulation: Interfaces in Computer Science and Operations Research, edited by Manuel Laguna and José Luis Gonzáles-Velarde. 2nd ed. 1–24. Boston, MA: Cambridge University Press.

This was produced by typing:

\begin{thebibliography}{12}
\bibitem[Agutter(1995)]{Agu95}
Agutter, A.~J. 1995. ``The Linguistic Significance of Current {B}ritish
 Slang.'' Unpublished doctoral dissertation, Edinburgh University, UK.
\bibitem[Barlow(1991)]{Bar91}
Barlow, D.~H. 1991. ``Diagnoses, Dimensions, and {DSM-IV} [Special
 issue].'' \emph{Journal of Abnormal Psychology} 100 (3).
\bibitem[Ericsson and Simon(1984)]{Eri1984}
Ericsson, K.~Anders, and Herbert~A. Simon. 1984. \emph{Protocol
 Analysis: Verbal Reports as Data}. Cambridge, MA: MIT Press.
\bibitem[Evans(1994)]{ev94}
Evans, W.~A. 1994. ``Approaches to Intelligent Information Retrieval.''
 \emph{Information Processing and Management} 7 (2): 147--168.
\bibitem[Glover(1986)]{glov86}
Glover, Fred. 1986. \emph{Hilbert Modular Forms and the {G}alois
 Representations Associated to {H}ilbert--{B}lumenthal Abelian
 Varieties}. Ph.D. thesis, Harvard University, Cambridge, MA.
\bibitem[Glover, Smith, and Ribeiro(1951)]{GloRib51}
Glover, Fred, Raymond Smith, and C.~C. Ribeiro, eds. 1951.
 \emph{Lessons of the {B}ritish War Economy}. vol.~5. 2nd ed.
 Westport, CT: Greenwood Press.
\bibitem[Holland(2004)]{Holl04}
Holland, M. 2004. ``Guide to Citing Internet Sources.'' Accessed
 November 4, 2012.
 http://www.bournemouth.ac.uk/library/using/guide\_to\_citing\_internet\_sourc.html.

\bibitem[Kern(1997)]{hk97}
Kern, Hudson. 1997. ``The Resurgent {J}apanese Economy and a
 {J}apan--{U}nited {S}tates Free Trade Agreement.'' In \emph{4th
 International Conference on the Restructuring of the Economic and
 Political System in {J}apan and {E}urope}, edited by C.~Lambert
 and G.~Holst. Milan, Italy, 21--25 May 1996. 147--156. Singapore:
 World Scientific.
\bibitem[Mazzeo(1991)]{Maz91}
Mazzeo, J. 1991. \emph{Comparability of Computer and Paper-and-Pencil
 Scores} (College Board Rep. No. 91). Princeton, NJ: Educational
 Testing Service.
\bibitem[Miller(1993)]{Mil93}
Miller, M.~E. 1993. \emph{The Interactive Tester (Version 4.0)
 [Computer software]}. Westminster, CA: Psytek Services.
\bibitem[Misner(1973)]{cwm73}
Misner, Charles~W. 1973. ``Efficient Algorithms for Layer Assignment
 Problems.'' In \emph{Gravitation in a Collapsing {U}niverse}, Vol.~5
 of \emph{Einstein's Legacy}, edited by Ignatius Gottlob. 2nd ed.
 63--83. San Francisco, CA: Freeman.
\bibitem[Pierce et~al.(1976)]{PeaEtAl76}
Pierce, Ian~F., Fred Glover, A.~T. Ogilvy, and A.~R.~S. Clements. 1976.
 ``A Model of Output, Employment, Wages and Prices in the {UK}.'' In
 \emph{Computing Tools for Modeling, Optimization and Simulation:
 Interfaces in Computer Science and Operations Research}, edited by
 Manuel Laguna and Jos\'{e}~Luis Gonz\'{a}les-Velarde. 2nd ed. 1--24.
 Boston, MA: Cambridge University Press.
\end{thebibliography}

Each entry takes the form:

\bibitem[authors' names(date of publication)]{key} Bibliography entry

where ‘authors' names’ is the list of names to appear where the bibitem is cited in the text, and ‘key’ is the tag that is to be used as an argument for the \cite{} commands in the text of the article. Bibliography entry should be the material that is to appear in the list of references, suitably formatted.

Instead of typing the bibliography by hand, you may prefer to create the list of references using a BibTEX database. Include the lines

\bibliographystyle{gENO}
\bibliography{gENOguide}

where the list of references should appear, where gENO.bst is the BibTEX style file for this journal and gENOguide.bib is the database of bibliographic details for the references section included with the gENO L^{A}TEX style guide package (to be replaced with the name of your own BibTEX database). The L^{A}TEX source file of your article will extract from

Table 2. gENO macros.

\thanks{title-page footnote to article title  e.g. ‘Corresponding author. E-mail:
or author}  A.N. Author@uiowa.edu’
\begin{abstract}...\end{abstract}  for abstract on titlepage
\bm{math and symbols}  bold italic math and symbols
\bi{text}  bold italic text
\sf{text or upright symbols in math mode} sans serif text or upright symbols in math mode

your .bib file only those references that are cited in that article and list them in the References section of it.

Please include either a copy of your .bib file or the final generated .bbl file among your source files if your .tex file does not contain a reference list in a thebibliography environment.

4.9 Appendices

Any appendices should be placed after the list of references, beginning with the command \appendices followed by the command \section for each appendix title, e.g.

\appendices
\section{This is the title of the first appendix}
\section{This is the title of the second appendix}

produces:

Appendix A. This is the title of the first appendix

Appendix B. This is the title of the second appendix

Subsections, equations, figures, tables, etc. within appendices will then be automatically numbered as appropriate.

4.10 gENO macros

Table 2 gives a list of macros for use with gENO. The list displays each macro’s code and a description/demonstration of its function.

5. Example of a section heading including small caps, italic, and bold Greek such as κ

The following code shows how to achieve this section head:

\section{Example of a section heading including
{\fontencoding{T1}\scshape{small caps}}, {\bi italic},
and bold Greek such as ${\bm\kappa}$}\label{headings}

6. gENO journal style

The notes given here relate to common style errors found in manuscripts, but are not intended to be exhaustive.

6.1 Hyphens, n-rules, m-rules and minus signs

(i) Hyphens (one dash in TEX/L^{A}TEX). gENO uses hyphens for compound adjectives (e.g. low-density gas, least-squares fit, two-component model) but not for complex units or ranges, which could become cumbersome (e.g. 15 km s^{−1} feature, 100–200 μm observations).

(ii) n-rules (two dashes in TEX/L^{A}TEX). These are used (a) to denote a range (e.g. 1.6– 2.2 μm); (b) to denote the joining of two words of equal standing (e.g. Kolmogorov– Smirnov test, Herbig–Haro object).

(iii) The m-rule (three dashes in TEX/L^{A}TEX) may be used as an alternative to parentheses

(e.g. ‘the results—assuming no temperature gradient—are indicative of … ’). (iv) The minus sign (one dash in TEX/L^{A}TEX) is produced automatically in math mode

by use of a single dash, e.g.

y_{i} ∈ {−1, 1} ∀i ∈ V,  (4)

where | − V | = A^{2} + B^{2}.

is produced by

   \begin{equation}
   y_{i} \in \{-1, 1 \} \quad \forall i \in V,
   \end{equation}
   \noindent where $|-V|=A^2+B^2.$

6.2 References

It is important to use the correct reference style, details of which can be found in Section 4.8 above.

6.3 Maths fonts

Scalar variables should be mediumface italic (e.g. s for speed); vectors should be bold italic (e.g. v for velocity); matrices should be bold roman (upright) (e.g. A), and tensors should be bold upright sans serif (e.g. L). Differential d, partial differential ∂, complex i, exponential e, superscript T for ‘transpose’, sin, cos, tan, log, etc., should all be roman. Openface, or ‘blackboard’, fonts can be used, for example, for the integers ℤ and the reals ℝ. Sub/superscripts that are physical variables should be italic, while those that are labels should be roman (e.g. C_{p}, T_{eff}). Displayed equations should have end-of-line punctuation appropriate to the running text sentence of which they form a part.

7. Troubleshooting

Authors may from time to time encounter problems with the preparation of their articles in L^{A}TEX. The appropriate action to take will depend on the nature of the problem—the following is intended to act as a guide.

(i) If the problem is with L^{A}TEX itself, rather than with the actual macros, please refer to the appropriate handbooks for initial advice.^{1} If the solution cannot be found,

1TEX: Knuth, D., 1986, The TEX Book (New York: Addison–Wesley); L^{A}TEX 2ε: Lamport, L., 1994, L^{A}TEX: A Document Preparation System, 2nd ed. (New York: Addison–Wesley).

and you suspect that the problem lies with the macros, then please contact Taylor

& Francis (latex.helpdesk@tandf.co.uk).

(ii) Problems with page make-up (e.g. large spaces between paragraphs, or under headings or figures; uneven columns; figures/tables appearing out of order): please do not attempt to remedy these yourself using ‘hard’ page make-up commands—the typesetter will correct such problems. (You may, if you wish, draw attention to particular problems when submitting the final version of your article.)

(iii) If a required font is not available at your site, allow TEX to substitute the font and specify which font you require in the covering letter accompanying your file(s).

8. Fixes for coding problems

This guide has been designed to minimize the need for user-defined macros to create special symbols. Authors are urged, wherever possible, to use the following coding rather than to create their own. This will minimize the danger of author-defined macros being accidentally ‘overridden’ when the article is typeset (see Section 4.5, ‘Typesetting mathematics’). In cases where it is essential to create your own macros, these should be displayed in the preamble of the source file before \begin{document}.

(i) Fonts in section headings and article titles. The following are examples of styles that sometimes prove difficult to code.

Article titles:

Generalized Flory theory at δ > 50^{∘}

is produced by

\title{Generalized Flory theory at
        ${\bm\delta > {\bfseries 50}^\circ}$}

Ion–ion correlations in H ii regions

is produced by

\title{Ion--ion correlations in H\,{\sc ii} regions}

(ii) n-rules, m-rules, hyphens and minus signs (see Section 6.1 for correct usage). To create the correct symbols in the sentence

The high-resolution observations were made along a line at an angle of −15^{∘} (East from North) from the axis of the jet—which runs North–South.

you would use the following code:

The high-resolution observations were made along a line at an
angle of $-15^\circ$ (East from North) from the axis of the
jet---which runs North--South.

(iii) Fonts in superscripts and subscripts. Subscripts and superscripts will automatically come out in the correct font and size in a math environment (e.g. enclosed by ‘$’ delimiters in running text or within \[...\] or the ‘equation’ environment for displayed equations). You can create the output k_{x} by typing ${\bm k_x}$. If the subscripts or superscripts need to be other than italic, they should be coded individually—see (vi) below.

(iv) Calligraphic letters (uppercase only). Normal calligraphic can be produced with \cal

as usual (in math mode).

(v) Automatic scaling of brackets. The codes \left and \right should be used to scale brackets automatically to fit the equation being set. For example, to get

v = x (\frac{N + 2}{N})

use the code

\[
   v = x \left( \frac{N+2}{N} \right)
\]

(vi) Roman font in equations. It is often necessary to make some symbols roman in an equation (e.g. units, non-variable subscripts). For example, to get

σ ≃ (r/13 h^{−1} Mpc)^{−0.9}, ω = \frac{N − N_{s}}{N_{R}}

use the code:

\[
   \sigma \simeq (r/13~h^{-1}
   ~{\rm Mpc})^{-0.9}, \qquad \omega
   =\frac{N-N_{{\rm s}}}{N_{{\rm R}}}
\]

The siunits package of macros for typesetting units is also compatible with the gENO class file.

9. Obtaining the gENO2e class file

9.1 Via the Taylor & Francis website

This Guide for Authors and the gENO2e.cls class file may be obtained via the Instructions for Authors on the Taylor & Francis homepage for the journal.

Please note that the class file calls up the following open-source L^{A}TEX packages, which will, for convenience, unpack with the downloaded Guide for Authors and class file: amsbsy.sty; amsfonts.sty; amsmath.sty; amssymb.sty; epsfig.sty; graphicx.sty; natbib.sty; rotating.sty; subfigure.sty.

9.2 Via e-mail

This Guide for Authors, the class file and the associated open-source L^{A}TEX packages are also available by e-mail. Requests should be addressed to latex.helpdesk@tandf.co.uk clearly stating for which journal you require the Guide for Authors and/or class file.